# A PNP ion channel deep learning solver with local neural network and finite element input data

### v1

@@ 1 Introduction
Ion channels are embedded within cell membranes, playing crucial roles in many important physiological processes including nerve signaling, muscle contraction, and cell communication [1]. As one important dielectric continuum model for simulating ion transport across biological membranes, a Poisson-Nernst-Planck ion channel (PNPic) model has been developed and widely used in biophysics and computational biology to calculate macroscopic ion channel kinetics (e.g. Gibbs free energy, electric currents, transport fluxes, membrane potential, and electrochemical potential) due to its significant advantages in computational efficiency. Hence, developing its effective numerical methods can have significant impacts on ion channel studying. Mathematically, the PNPic model is a system of nonlinear partial differential equations (PDE) for computing ionic concentrations and an electrostatic potential function. It has been numerically solved by using finite difference [2, 3], finite element [4, 5, 6, 7, 8], and integral equation methods [9]. In ion channel study and simulation, it is often required to solve the PNPic model numerically on a large mesh domain with a minuscule mesh size to explore the properties of an ion channel protein in terms of ionic concentration and electrostatic potential functions since these functions may change intensely within an ion channel pore and areas near the ion channel protein and membrane regions. However, doing so is still very expensive in numerical calculations. On the other hand, during an ion channel study, some model parameters, interface values, and boundary values are often required to adjust properly. With a traditional numerical partial differential equation method (e.g., the finite element method), we have to resolve the model for each change of the parameters, further increasing the simulation study costs.
@@
Several PDE deep learning methods have been reported in the literature [10, 11, 12, 13]. However, they rely on a global neural network, making them have higher complexities in calculation and implementation, and demanding a large number of fine grid simulations for training data. To sharply reduce their complexities, recently, a neural network with local converging inputs (NNLCI) is introduced in [14, 15], opening a promising way for developing machine learning algorithms for solving 1D, 2D, and 3D PDE problems. It has been shown to work for 1D and 2D Euler’s equations and electromagnetic equations, which may contain shocks, contact discontinuities, and their interactions for gas dynamics. It also has been successfully applied to predicting electromagnetic waves scattered around complicated perfect electric conductors in [16]. In [17], NNLCI has been developed for predicting supersonic flows in irregular domains with unstructured grids. NNLCI lets a trained neural network function like a local scanner which scans two coarse grid numerical solutions (with one more accurate than the other) pointwisely and then predicts the high-fidelity solution values at each scanned location. Therefore the neural network is local - taking input from local patches of the numerical solutions and generating output as the high-fidelity solution values at the center of the patches. The localness provides many nice features of the approach. Here we only list five of them as follows:
@@
(I) The local neural network can be small and simple. (II) The demand for training data is minimal because a fine-grid simulation can provide hundreds or thousands of local samples for training, and a couple of or tens of fine-grid simulations are usually adequate during the training process. (III) The fine-grid simulations for training can be placed sparsely (thus covering more variations of the problem) since the local samples cropped from different solutions may still bear some similarities. (IV) It can predict solutions containing discontinuities (such as shocks) sharply. (V) The training can be done on a set of computational domains and the prediction can be done on different domains.
@@
However, it’s uncertain whether or not NNLCI will work beyond hyperbolic systems. In this work, we develop a PNPic deep learning solver to enable us to get a PNPic numerical solution in high accuracy efficiently for a range of parameters, interface values, and boundary values. This work is our first attempt to extend NNLCI to a very different PDE system with convection-diffusion equations and elliptic equations. It is also the first time for us to adapt NNLCI to deal with multiple subdomains connected with complicated interface conditions, with a distinct system of equations defined in each subdomain. With this new solver, we can address two fundamental questions in a simple and elegant manor: Does NNLCI work for elliptic systems given that prior arts are for hyperbolic systems? And how to deal with multi subdomains (and different nonlinear elliptic systems) with complicated interface conditions efficiently? The simplicity of our solver bodes well for our future extension to a complicated problem involving a 3D molecular structure of an ion channel protein. The 1D PNPic model is also one important model by itself in biophysics since it has been validated and applied to the simulation of ionic transport and the studies of various membrane kinetics such as membrane potentials, conductance, transport fluxes, electric current, etc. Hence, our 1D PNPic deep learning solver is valuable in biophysics, having an important impact on ion channel studies.
@@
In particular, our PNPic deep learning solver is defined as a combination of a new 1D PNPic finite element solver with a local neural network. We adapt NNLCI here to reduce the complexity of finite element schemes when repeated simulations with varying setups of the model are conducted. We demonstrate that NNLCI works not only for hyperbolic equations as in prior works but also for the nonlinear PDE systems in multiple subdomains connected with complicated interface conditions. This is a simple feed forward neural network which takes as input local patches of approximated solutions on two low resolution meshes, one with a slightly higher resolution. The neural network is able to model the local dynamics between the solutions on the two different resolutions to predict an even more accurate solution. Thus once trained, the model is able to transform low-cost solutions into a high fidelity solution, by piecing together all the local predictions. The biggest cost is to generate reference solutions with the initial finite element method for training. However, compared with conventional global neural network methods [10, 11, 12, 13], our PNPic solver has a much lower dimensional input space as it only predicts on local patches. Thus, it can generate enough training data from a few training data to be accurately predictive on a broad range of very different solutions in high accuracy. Moreover, as a new numerical method for solving a nonlinear interface boundary value problem, the novel wisdom is to assign a neural network for each subdomain and address the messy interface conditions between these subdomains. We take a bold approach — use a single neural network to cover all subdomains, with no need to deal with any interface conditions! This simplicity of the methodology itself implies a fundamental breakthrough for solving these types of complicated problems with great robustness and efficiency for practical use.
@@
In this paper, we first revisit the 1D PNPic model to redefine it as a system of nonlinear variational problems. We then develop a finite element iterative algorithm for solving our 1D model and implement it as a software package in Python based on the state-of-the-art finite element library from the FEniCS project [18], resulting in an efficient PNPic finite element solver, which is used to generate data for training and testing our local neural network as well as the PNPic reference solution required in a network training process. We next define the 1D PNPic deep learning solver as a combination of our finite element solver with a local neural network scheme such that a 1D PNPic numerical solution with a high accuracy close to the one defined in a very fine grid can be generated by using a local neural network and a set of coarse grid numerical solutions, which has a low accuracy but can be cheaply generated by a PNPic finite element solver. Remarkably, our local neural network can be trained much faster than a conventional global neural network since its input data only involves a small number of local coarse numerical solutions. Furthermore, the local neural network can work for different models disturbed by changing the parameters within a proper range provided that it is trained properly by using a database that collects numerical solutions reflecting different adjustment cases. Numerical results are finally reported, demonstrating that the PNPic deep learning solver can efficiently generate numerical solutions with a high solution accuracy.
@@ 2.1 The Poisson-Nernst-Planck ion channel model
We consider a PNPic model for a potassium (${}^{+}$) channel and a solvent with two ion species in the steady state. Following the modeling work done in [19, 9], we split the channel pore domain into three regions — an interior bath region, a channel region, and an exterior bath region. The two baths are set as cones while the channel is set as a cylinder. To reflect the biochemical features of a potassium channel in the model, we further split the channel region into four subdomains, called the buffer, nonpolar, central cavity, and selectivity filter, respectively. Because of the rotational symmetry, we can set the $x$-axis to go through the center of the channel pore such that the channel pore domain can be described as an interval, $[a,b]$, and the six subdomains can be represented by the six sub-intervals of the interval $[a,b]$ as illustrated in Fig. 1 with the five partition numbers $x_{i}$ for $i=1,2,3,4,5$. The cross-sectional area $A$ of the channel pore can then be calculated by the formula $A(x)=\pi r(x)^{2}$ for $a\leq x\leq b$ with $r$ being a radius. Here $r$ is a constant in the channel region interval $[x_{1},x_{5}]$, a linear function of $x$ in the interior bath region interval $[a,x_{1}]$, and another linear function in the exterior bath region interval $[x_{5},b]$.
Figure 1: A cross section of the ${}^{+}$ ion channel model used in [19, 9]. Here the channel pore domain consists of the interior and exterior conical baths and the cylindrical channel while the channel is further split into the buffer, nonpolar, central cavity, and selectivity filter subdomains to characterize its biochemical properties.
@@
-- $A(x)^{-1}\frac{d}{dx}\left[\epsilon(x)A(x)\frac{d\phi(x)}{dx}\right]=\frac{e_{c}}{\epsilon_{0}}\left[z_{1}c_{1}(x)+z_{2}c_{2}(x)+\rho(x)\right],\quad a<x<b,$ (1a)
$A(x)^{-1}\frac{d}{dx}\left[A(x)D_{i}(x)\left(\frac{dc_{i}(x)}{dx}+z_{i}\frac{e_{c}}{k_{B}T}c_{i}(x)\frac{d\phi(x)}{dx}\right)\right]=0,\quad i=1,2,\quad a<x<b,$ (1b)
subject to Dirichlet boundary conditions
@@
where $\epsilon_{0}$ is the permittivity constant of vacuum, $\epsilon$ is the permittivity function, $z_{1}$ and $z_{2}$ are the charge numbers for anions and cations, respectively, $D_{i}$ is a diffusion function of species $i$, $k_{B}$ is the Boltzmann constant, $T$ is the absolute temperature, $e_{c}$ is the elementary charge, $\phi_{a},\phi_{b}$, $c_{i,a}$, and $c_{i,b}$ are given boundary values, and $\rho$ is a permanent charge density, which is used to account for the charge effect from an ion channel protein. Here $\epsilon$ and $D_{i}$ are differentiable and the units of $e_{c},T$, $k_{B}$, $\phi$, and $c_{i}$ are in the International System of Units. Thus, $k_{B}T/e_{c}$ is about 0.026 volts at $T=298.15$ Kelvin. Eq. 1a is the Poisson equation, and the two equations of Eq. 1b are the Nernst-Planck equations.
@@
where $\epsilon_{ib},\epsilon_{b},\epsilon_{n},\epsilon_{c},\epsilon_{f},\epsilon_{xb},$ $D_{ib},D_{b},D_{n},D_{c},D_{f}$, and $D_{xb}$ are positive numbers. In this case, we have to formulate the PNPic model defined in Eq. 2.1 as an interface boundary value problem, which involves six equations defined in the six sub-intervals and connected by interface conditions. We do not present this formulation instead of reformulating Eq. 2.1 into a variational problem since we solve the PNP ion channel model by a finite element method in this work.
Multiplying each equation of Eqs. 1a and  1b by a test function, $v$, and using integration by parts, we can get the variational problem as follows: Find $\phi\in H^{1}(a,b)$ and $c_{i}\in H^{1}(a,b)$ for $i=1,2$ satisfying the Dirichlet boundary value conditions of Eq. 2.1 such that
$\int_{a}^{b}\epsilon(x)A(x)\frac{d\phi(x)}{dx}\frac{dv(x)}{dx}\mathrm{~{}d}x$displaystyle-- $\frac{e_{c}}{\epsilon_{0}}\int_{a}^{b}A(x)[z_{1}c_{1}(x)+z_{2}c_{2}(x)]v(x)\mathrm{~{}d}x$ (4)
$=$ $\frac{e_{c}}{\epsilon_{0}}\int_{a}^{b}A(x)\rho(x)v(x)\mathrm{~{}d}x\quad\forall v\in H_{0}^{1}(a,b),$
$\int_{a}^{b}A(x)D_{i}(x)\left[\frac{dc_{i}(x)}{dx}+z_{i}\frac{e_{c}}{k_{B}T}c_{i}(x)\frac{d\phi(x)}{dx}\right]\frac{dv_{i}(x)}{dx}\mathrm{~{}d}x=0,\quad i=1,2,\quad\forall v\in H_{0}^{1}(a,b),$ (5)
where $H^{1}(a,b)$ denotes a Sobolev function space in which all the functions have the first derivatives over the interval $(a,b)$ and $H_{0}^{1}(a,b)=\{v\in H^{1}(a,b)\mid v(a)=0,v(b)=0\}$.
@@ 2.2 The PNPic finite element iterative solver
Substituting $V$ and $V_{0}$ to $H^{1}(a,b)$ and $H_{0}^{1}(a,b)$, respectively, we immediately derive a system of finite element equations from the variational equations of Eqs. 4 and 5. We propose to solve this system by a damped iterative scheme as defined below.
Let $\phi^{(k)}$ and $c_{i}^{(k)}$ denote the $k$-th iterates of $\phi$ and $c_{i}$, respectively. When the initial iterates $\phi^{(0)}$ and $c_{i}^{(0)}$ are given, for $k=0,1,2,\ldots$, we define the updates $\phi^{(k+1)}$ and $c_{i}^{(k+1)}$ by
$c_{i}^{(k+1)}=c_{i}^{(k)}+\omega\left(p_{i}-c_{i}^{(k)}\right),\quad i=1,2,$
@@
where $\omega$ is a damping parameter between 0 and 1, $p_{i}$ is a solution of the linear finite element variational problem: Find $p_{i}\in V$ satisfying $p_{i}(a)=c_{i,a}$ and $p_{i}(b)=c_{i,b}$ for $i=1,2$ such that
$\int_{a}^{b}A(x)D_{i}(x)\left[\frac{dp_{i}(x)}{dx}+z_{i}\frac{e_{c}}{k_{B}T}p_{i}(x)\frac{d\phi^{(k)}(x)}{dx}\right]\frac{dv_{i}(x)}{dx}\mathrm{~{}d}x=0,\quad i=1,2,\quad\forall v_{i}\in V_{0},$ (6)
@@
$\int_{a}^{b}\epsilon(x)A(x)\frac{dq(x)}{dx}\frac{dv(x)}{dx}\mathrm{~{}d}x$ $=$ $\frac{e_{c}}{\epsilon_{0}}\int_{a}^{b}A(x)[z_{1}c_{1}^{(k+1)}(x)+z_{2}c_{2}^{(k+1)}(x)]v(x)\mathrm{~{}d}x$ (7)
$-\frac{e_{c}}{\epsilon_{0}}\int_{a}^{b}A(x)\rho(x)v(x)\mathrm{~{}d}x\quad\forall v\in V_{0},$
@@
We terminate the iterative process whenever the following iteration termination rule holds:
@@
We implemented the above finite element iterative solver as a Python software package based on the state-of-the-art finite element library from the FEniCS project [18]. We then validated this package numerically by using an analytical solution of the PNPic model Eq. 2.1. See Appendix for details.

### v2

@@ 1 Introduction
Ion channels, embedded within cell membranes, play crucial roles in many important physiological processes including nerve signaling, muscle contraction, and cell communication [10]. As one important dielectric continuum model for simulating ion transport across biological membranes, a Poisson-Nernst-Planck ion channel (PNPic) model has been developed and widely used in biophysics and computational biology to calculate macroscopic ion channel kinetics (e.g., Gibbs free energy, electric currents, transport fluxes, membrane potential, and electrochemical potential) due to its significant advantages in computational efficiency. Hence, developing its effective numerical methods can have significant impacts on ion channel studying. Mathematically, the PNPic model is a system of nonlinear partial differential equations (PDE) for computing ionic concentrations and an electrostatic potential function. Various numerical methods including finite difference [3, 23], finite element [24, 21, 5, 20, 22], and integral equation methods [4] have been used to solve this model. In ion channel study and simulation, it is often required to solve the PNPic model numerically on a large mesh domain with a minuscule mesh size to explore the properties of an ion channel protein in terms of ionic concentration and electrostatic potential functions since these functions may change intensely within an ion channel pore and areas near the ion channel protein and membrane regions. However, doing so is still very expensive in numerical calculations. On the other hand, during an ion channel study, some model parameters, interface values, and boundary values are often required to adjust properly. With a traditional numerical partial differential equation method (e.g., the finite element method), we have to resolve the model for each change of the parameters, further increasing the simulation study costs.
@@
Several PDE deep learning methods have been reported in the literature [6, 14, 13, 16, 1]. However, they rely on a global neural network, making them have higher complexities in calculation and implementation, and demanding a large number of fine grid simulations for training data. To sharply reduce their complexities, recently, a neural network with local converging inputs (NNLCI) has been introduced in [11, 12], opening a promising way for developing machine learning algorithms for solving one-dimensional (1D), two-dimensional (2D), and three-dimensional (3D) PDE problems. It has been shown to work for 1D and 2D Euler’s equations and electromagnetic equations, which may contain shocks, contact discontinuities, and their interactions for gas dynamics. It also has been successfully applied to predicting electromagnetic waves scattered around complicated perfect electric conductors in [7]. In [8], NNLCI has been developed for predicting supersonic flows in irregular domains with unstructured grids.
@@
NNLCI lets a trained neural network function like a local scanner which scans two coarse grid numerical solutions (with one more accurate than the other) point-wisely and then predicts the high-fidelity solution values at each scanned location. Therefore the neural network is local - taking input from local patches of the numerical solutions and generating output as the high-fidelity solution values at the center of the patches. The localness provides many nice features of the approach. Here we only list five of them: (I) The local neural network can be small and simple. (II) The demand for training data is minimal because a fine-grid simulation can provide hundreds or thousands of local samples for training, and a couple of or tens of fine-grid simulations are usually adequate during the training process. (III) The fine-grid simulations for training can be placed sparsely (thus covering more variations of the problem) since the local samples cropped from different solutions may still bear some similarities. (IV) It can predict solutions containing discontinuities (such as shocks) sharply. (V) The training can be done on a set of computational domains and the prediction can be done on different domains.
@@
However, it’s uncertain whether or not NNLCI will work beyond hyperbolic systems. In this work, we develop deep learning techniques to enable us to get a PNPic numerical solution in high accuracy efficiently for a range of parameters, interface values, and boundary values. This work is our first attempt to extend NNLCI to a very different PDE system with convection-diffusion equations and elliptic equations. It is also the first time for us to adapt NNLCI to deal with multiple subdomains connected with complicated interface conditions, with a distinct system of equations defined in each subdomain. With a deep learning solver, we can address two fundamental questions in a simple and elegant manor: Does NNLCI work for elliptic systems given that prior arts are for hyperbolic systems? And how to deal with multi subdomains (and different nonlinear elliptic systems) with complicated interface conditions efficiently?
As the first step toward the direction of developing PNPic deep learning solvers, in this work, we develop a PNPic deep learning method for solving a 1D PNPic model proposed in [9]. In the current literature, the majority of PNPic models for potassium channels are constructed in 1D or quasi-1D geometries. By reducing the 3D dimension to a single coordinate representing the axis along the channel pore, 1D models become computationally tractable while still maintaining basic ion channel properties. After being validated by experimental data, they are often used by biochemists, biophysicists, and bioengineers in their study of ion transport phenomena in channels. The 1D PNPic model that we select for developing PNPic deep learning methods is one important model by itself. It has been validated and applied to the simulation of ionic transport and the studies of various membrane kinetics such as membrane potentials, conductance, transport fluxes, electric current, etc [19, 17, 2, 18]. It was recently selected to develop a boundary element method in [4]. Hence, our 1D PNPic deep learning solver is valuable in biophysics, having an important impact on ion channel studies. Moreover, the simplicity of our solver bodes well for our future extension to a complicated problem involving a 3D molecular structure of an ion channel protein.
@@
In particular, our PNPic deep learning solver is defined as a combination of a new 1D PNPic finite element solver with a local neural network. We adapt NNLCI here to reduce the complexity of finite element schemes when repeated simulations with varying setups of the model are conducted. We demonstrate that NNLCI works not only for hyperbolic equations as in prior works but also for the nonlinear PDE systems in multiple subdomains connected with complicated interface conditions. This is a simple feed-forward neural network that takes as input local patches of approximated solutions on two low-resolution meshes, one with a slightly higher resolution. The neural network can model the local dynamics between the solutions on the two different resolutions to predict an even more accurate solution. Thus once trained, the model can transform low-cost solutions into a high fidelity solution, by piecing together all the local predictions. The biggest cost is to generate reference solutions with the initial finite element method for training. However, compared with conventional global neural network methods [6, 14, 13, 16], our PNPic solver has a much lower dimensional input space as it only predicts on local patches. Thus, it can generate enough training data from a few training data to be accurately predictive on a broad range of very different solutions in high accuracy. Moreover, as a new numerical method for solving a nonlinear interface boundary value problem, the novel wisdom is to assign a neural network for each subdomain and address the messy interface conditions between these subdomains. We take a bold approach — use a single neural network to cover all subdomains, with no need to deal with any interface conditions! This simplicity of the methodology itself implies a fundamental breakthrough for solving these types of complicated problems with great robustness and efficiency for practical use.
@@
In this paper, we first revisit the 1D PNPic model and redefine it as a system of nonlinear variational problems. We then develop a finite element iterative algorithm for solving our 1D model and implement it as a software package in Python based on the state-of-the-art finite element library from the FEniCS project [15], resulting in an efficient PNPic finite element solver, which is used to generate data for training and testing our local neural network as well as the PNPic reference solution required in a network training process. We next define the 1D PNPic deep learning solver as a combination of our finite element solver with a local neural network scheme such that a 1D PNPic numerical solution with a high accuracy close to the one defined in a very fine grid can be generated by using a local neural network and a set of coarse grid numerical solutions, which has a low accuracy but can be cheaply generated by a PNPic finite element solver. Remarkably, our local neural network can be trained much faster than a conventional global neural network since its input data only involves a small number of local coarse numerical solutions. Furthermore, the local neural network can work for different models disturbed by changing the parameters within a proper range provided that it is trained properly by using a database that collects numerical solutions reflecting different adjustment cases. Numerical results are finally reported, demonstrating that the PNPic deep learning solver can efficiently generate numerical solutions with a high solution accuracy.
@@ 2.1 The Poisson-Nernst-Planck ion channel model
We consider a 1D PNPic model for a potassium (K+) channel and a solvent with two ion species in the steady state. This model is first proposed in [9] and has been validated by experimental data and subsequent research that cites or builds upon that work. As needed to develop a finite element method for solving this model, we revisit it in our notation and reformulate it in a variational form. In particular, we split the channel pore domain into three regions — an interior bath region, a channel region, and an exterior bath region. The two baths are set as cones while the channel is set as a cylinder. To reflect the biochemical features of a potassium channel in the model, we further split the channel region into four subdomains, called the buffer, nonpolar, central cavity, and selectivity filter, respectively. Because of the rotational symmetry, we can set the $x$-axis to go through the center of the channel pore such that the channel pore domain can be described as an interval, $[a,b]$, and the six subdomains can be represented by the six sub-intervals of the interval $[a,b]$ as illustrated in Figure 1 with the five partition numbers $x_{i}$ for $i=1,2,3,4,5$. The cross-sectional area $A$ of the channel pore can then be calculated by the formula $A(x)=\pi r(x)^{2}$ for $a\leq x\leq b$ with $r$ being a radius. Here $r$ is a constant in the channel region interval $[x_{1},x_{5}]$, a linear function of $x$ in the interior bath region interval $[a,x_{1}]$, and another linear function in the exterior bath region interval $[x_{5},b]$.
Figure 1: A cross section of the K+ ion channel model used in [9, 4]. Here the channel pore domain consists of the interior and exterior conical baths and the cylindrical channel while the channel is further split into the buffer, nonpolar, central cavity, and selectivity filter subdomains to characterize its biochemical properties.
@@
-- $A(x)^{-1}\frac{d}{dx}\left[\epsilon_{0}\epsilon(x)A(x)\frac{d\phi(x)}{dx}\right]={e_{c}}\left[z_{1}c_{1}(x)+z_{2}c_{2}(x)\right]+\rho(x),\quad a<x<b,$ (1a)
$A(x)^{-1}\frac{d}{dx}\left[A(x)D_{i}(x)\left(\frac{dc_{i}(x)}{dx}+z_{i}\frac{e_{c}}{k_{B}T}c_{i}(x)\frac{d\phi(x)}{dx}\right)\right]=0,\quad a<x<b,$ (1b)
for $i=1,2$, along with Dirichlet boundary conditions
@@
where $\epsilon_{0}$ is the permittivity constant of vacuum, $\epsilon$ is the permittivity function, $z_{1}$ and $z_{2}$ are the charge numbers for anions and cations, respectively, $D_{i}$ is a diffusion function of species $i$, $k_{B}$ is the Boltzmann constant, $T$ is the absolute temperature, $e_{c}$ is the elementary charge, $\phi_{a},\phi_{b}$, $c_{i,a}$, and $c_{i,b}$ are given boundary values, and $\rho$ is a permanent charge density, which is used to account for the charge effect from an ion channel protein. Here $\epsilon$ and $D_{i}$ are differentiable and the model is measured in the International System of Units. Thus, the units of $\phi$, $c_{i}$, $A(x)$, and $D_{i}$ are volts, per cube meters, square meters, and square meters per second, respectively, while $e_{c}/\epsilon_{0}$ is about $1.8095\times 10^{-8}$ volts meters and $k_{B}T/e_{c}$ is about 0.026 volts at $T=298.15$ Kelvin. Equation (1a) is the Poisson equation and the two equations of (1b) are the Nernst-Planck equations.
@@
where $\epsilon_{ib},\epsilon_{b},\epsilon_{n},\epsilon_{c},\epsilon_{f},\epsilon_{xb},$ $D_{ib},D_{b},D_{n},D_{c},D_{f}$, and $D_{xb}$ are positive numbers. In this case, we have to formulate the PNPic model defined in  (1) as an interface boundary value problem, which involves six equations defined in the six sub-intervals and connected by interface conditions. We do not present this formulation instead of reformulating  (1) into a variational problem since we solve the PNP ion channel model by a finite element method in this work.
Multiplying each equation of (1a) and  (1b) by a test function, $v$, and using integration by parts, we can get the variational problem as follows:
Find $\phi\in H^{1}(a,b)$ and $c_{i}\in H^{1}(a,b)$ for $i=1,2$ satisfying the Dirichlet boundary value conditions of  (2) such that
$\int_{a}^{b}\epsilon(x)A(x)\frac{d\phi(x)}{dx}\frac{dv(x)}{dx}\mathrm{~{}d}x-\frac{e_{c}}{\epsilon_{0}}\int_{a}^{b}A(x)\left[z_{1}c_{1}(x)+z_{2}c_{2}(x)\right]v(x)\mathrm{~{}d}x$ (4)
$=$ $\frac{1}{\epsilon_{0}}\int_{a}^{b}A(x)\rho(x)v(x)\mathrm{~{}d}x\quad\forall v\in H_{0}^{1}(a,b),$
$\int_{a}^{b}A(x)D_{i}(x)\left[\frac{dc_{i}(x)}{dx}+z_{i}\frac{e_{c}}{k_{B}T}c_{i}(x)\frac{d\phi(x)}{dx}\right]\frac{dv(x)}{dx}\mathrm{~{}d}x=0\quad\forall v\in H_{0}^{1}(a,b),$ (5)
for $i=1,2$. Here $H^{1}(a,b)$ is the Sobolev function space, in which all the functions have the first weak derivatives over the interval $(a,b)$, and $H_{0}^{1}(a,b)$ is defined by
$H_{0}^{1}(a,b)=\{v\in H^{1}(a,b)\mid v(a)=0,v(b)=0\}.$
@@ 2.2 The PNPic finite element iterative solver
Substituting $V$ and $V_{0}$ to $H^{1}(a,b)$ and $H_{0}^{1}(a,b)$, respectively, we immediately derive a system of finite element equations from the variational equations of (4) and (5). We propose to solve this system by a damped iterative scheme as defined below.
Let $\phi^{(k)}$ and $c_{i}^{(k)}$ denote the $k$-th iterates of $\phi$ and $c_{i}$, respectively. When the initial iterates $\phi^{(0)}$ and $c_{i}^{(0)}$ are given, we define the updates $\phi^{(k+1)}$ and $c_{i}^{(k+1)}$ by
$c_{i}^{(k+1)}=c_{i}^{(k)}+\omega\left(p_{i}-c_{i}^{(k)}\right),\quad i=1,2,$ (6)
@@
where $\omega$ is a damping parameter between 0 and 1, $p_{i}$ is a solution of the linear finite element variational problem: Find $p_{i}\in V$ satisfying $p_{i}(a)=c_{i,a}$ and $p_{i}(b)=c_{i,b}$ for $i=1,2$ such that for $i=1,2,$
$\int_{a}^{b}A(x)D_{i}(x)\left[\frac{dp_{i}(x)}{dx}+z_{i}\frac{e_{c}}{k_{B}T}p_{i}(x)\frac{d\phi^{(k)}(x)}{dx}\right]\frac{dv(x)}{dx}\mathrm{~{}d}x=0\quad\forall v\in V_{0},$ (7)
@@
$\int_{a}^{b}\epsilon(x)A(x)\frac{dq(x)}{dx}\frac{dv(x)}{dx}\mathrm{~{}d}x$ $=\frac{e_{c}}{\epsilon_{0}}\int_{a}^{b}A(x)\left[z_{1}c_{1}^{(k+1)}(x)+z_{2}c_{2}^{(k+1)}(x)\right]v(x)\mathrm{~{}d}x$ (8)
$\quad+\frac{1}{\epsilon_{0}}\int_{a}^{b}A(x)\rho(x)v(x)\mathrm{~{}d}x\quad\forall v\in V_{0},$
@@
We continue to carry out the iterative process (6) for $k=0,1,2,\ldots$ until the iteration termination rule holds:
@@
We implemented the above finite element iterative solver as a Python software package based on the state-of-the-art finite element library from the FEniCS project [15]. In the implementation, we scaled $\phi$ as a dimensionless potential function by the formula $\frac{e_{c}}{k_{B}T}\phi$. Following [9], we used nanometers, moles per liter, and $10^{-5}$ cm2 per second as the units of length, concentration, and diffusion, respectively. We also used the approximation values of some related physical constants (e.g. $e_{c}/(k_{B}T)$) given in [9]. In this way, we could produce the same test results as those reported in [9]. Furthermore, we validated our PNPic solver and package by using an analytical solution of the PNPic model (1). See Appendix for details.